Can someone help me solve this problem?
I need to multiply:
(-6) in the power of 2 - 4 in the power of 2
Thanks,
Lyn

Answer by fractalier(2101)

Sure...
(-6)^2 - 4^2 =
36 - 16 =
20
remember PEMDAS for your order of operations...